OpenStudy (anonymous):

Need help: 2a^2 + 30a + 100 Factoring!

hartnn (hartnn):

first factor out 2 , what u get ?

OpenStudy (anonymous):

\[2a ^{2} + 30a + 100\]

OpenStudy (anonymous):

That's what I find confusing about theis problem the factoring out of 2 . . . .

OpenStudy (anonymous):

Would it be \[a ^{2} + 15a +50\]?

hartnn (hartnn):

example : 2a+2cd+2 = 2 (a+cd+1) 2x^2+4x+8 = 2 (x^2 +2x+4)

OpenStudy (anonymous):

Oh so it would be \[2 (a ^{2} + a + 50)\]Roght?

OpenStudy (anonymous):

*Right

hartnn (hartnn):

did you mean 2 (a^2+15a+50) ??

OpenStudy (anonymous):

Yes *facepalm*

OpenStudy (anonymous):

So the I would have to find two numbers that add up to 12 but x's to 50?

OpenStudy (anonymous):

*then

hartnn (hartnn):

add up to 15 but x's to 50 yes, which are ?

OpenStudy (anonymous):

10 & 5

hartnn (hartnn):

yes, correct. so, 15a =10a+5a can you proceed ?

OpenStudy (anonymous):

Yes it would be: 2a^2 + 30a + 100 2 (a ^2 + 15a + 50) 10*5 = 50 10 + 5 =15 2 (a + 5) (a + 10)

hartnn (hartnn):

that.is.correct.

OpenStudy (anonymous):

That would finish the problem correct?

hartnn (hartnn):

absolutely :)

OpenStudy (anonymous):

Thanks . . . again :)

hartnn (hartnn):

welcome ... again :)